one number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by pwac(253)

Let x = smaller number
So,larger number = x + 12
Need to sort out an equation from information given
Smaller + twice the larger = 39. This gives us:
x + 2(x+12)= 39. Multiply expression in brackets by 2
x + 2x + 24 = 39 Add x terms
3x+24 = 39 Subtract 24 from both sides
3x = 15 Divide both sides by 3
x = 5 This is your smaller number
So, larger number is 5 + 12 = 17
Pete